OpenStudy (anonymous):

If 1cm=10km what is 36km and 41km?

OpenStudy (anonymous):

1cm/10km=x cm/36 km (1/10)*36 = x

OpenStudy (anonymous):

so 36 x 1/10?

OpenStudy (ajprincess):

ya.

OpenStudy (anonymous):

3 and 3 5s

OpenStudy (anonymous):

but with a ruler on paper?

OpenStudy (ajprincess):

I thnk u hav made a slight mistake @TheViper. 1cm=10km xcm=36km x=36/10 x=3.6 Can u do the other one?

OpenStudy (theviper):

yeah I was confused so i asked :)

OpenStudy (ajprincess):

Oh k. Nw do u get t?

OpenStudy (anonymous):

so basically 3.6cm and 4.1cm?

OpenStudy (theviper):

yes thanx @ajprincess :)

OpenStudy (anonymous):

Thanks guys for helping me! =D

OpenStudy (ajprincess):

ya right. @navpreet and u r welcome @TheViper.:)

OpenStudy (theviper):

:)

OpenStudy (anonymous):

thanks @ajprincess and @TheViper

OpenStudy (theviper):

me lol :) I was too confused :D Thanx @ajprincess :)

OpenStudy (ajprincess):

Welcome both of u.:)

mathslover (mathslover):

@navpreet : \[\huge{\mathcal{WELCOME}\textbf{TO}\mathbb{OPENSTUDY}}\]

OpenStudy (theviper):

Ambassador duty must :D

OpenStudy (theviper):

Nice \(LaTeX\)

mathslover (mathslover):

well right @TheViper

OpenStudy (anonymous):

thanks =D

mathslover (mathslover):

welcome navpreet.. though if you have any problem in this site you can message me.

OpenStudy (anonymous):

ok. :)

OpenStudy (theviper):

;)